OpenStudy (anonymous):

I HAVE 7 GEOMETRY QUESTIONS ANYONE THINK THEY CAN HELP?

OpenStudy (anonymous):

a. All of the examples in the above paragraph refer to integers that include only the digit 1. What would the integer 500 equal using powers of 10? How did you come up with that answer? (How do you use powers of 10 to get 500? I don't understand so an explanation would be greatly appreciated)

OpenStudy (anonymous):

Do you mean what power of 10 will give you 500 ? If so do 500 = 10^n then use ln and your n value will be approximately 2.7 and that will give you 500 back use 10?

OpenStudy (anonymous):

To be more mathematically correct n = ln(500)/ln(10) which will give you 500 back when you do 10^n.

OpenStudy (anonymous):

i think what she's actually saying is 500=n^10